\begin{thebibliography}{-------}
\providecommand{\natexlab}[1]{#1}

\bibitem[Feynman(1996)]{Feynman:1996kb}
Feynman, R.P.
\newblock {\em {Feynman lectures on gravitation}}; Reading, USA:
  Addison-Wesley,  1996.

\bibitem[Grignani and Nardelli(1992)]{Grignani:1991nj}
Grignani, G.; Nardelli, G.
\newblock {Gravity and the Poincare group}.
\newblock {\em Phys. Rev.} {\bf 1992}, {\em D45},~2719--2731.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1103/PhysRevD.45.2719}{\detokenize{10.1103/PhysRevD.45.2719}}}.

\bibitem[Cartan(1986)]{cartan1986manifolds}
Cartan, {\'E}.
\newblock {\em On Manifolds with an Affine Connection and the Theory of General
  Relativity}; Monographs and Textbooks in Physical Science, Bibliopolis,
  1986.

\bibitem[Aldrovandi and Pereira(2013)]{Aldrovandi:2013wha}
Aldrovandi, R.; Pereira, J.G.
\newblock {\em {Teleparallel Gravity}}; Vol. 173, Springer: Dordrecht,  2013.

\bibitem[Złośnik \em{et~al.}(2018)Złośnik, Urban, Marzola, and
  Koivisto]{Zlosnik:2018qvg}
Złośnik, T.; Urban, F.; Marzola, L.; Koivisto, T.
\newblock {Spacetime and dark matter from spontaneous breaking of Lorentz
  symmetry}.
\newblock {\em Class. Quant. Grav.} {\bf 2018}, {\em 35},~235003,
  \href{http://xxx.lanl.gov/abs/1807.01100}{{\normalfont
  [arXiv:gr-qc/1807.01100]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1088/1361-6382/aaea96}{\detokenize{10.1088/1361-6382/aaea96}}}.

\bibitem[Hehl \em{et~al.}(1995)Hehl, McCrea, Mielke, and Ne'eman]{Hehl:1994ue}
Hehl, F.W.; McCrea, J.D.; Mielke, E.W.; Ne'eman, Y.
\newblock {Metric affine gauge theory of gravity: Field equations, Noether
  identities, world spinors, and breaking of dilation invariance}.
\newblock {\em Phys. Rept.} {\bf 1995}, {\em 258},~1--171,
  \href{http://xxx.lanl.gov/abs/gr-qc/9402012}{{\normalfont
  [arXiv:gr-qc/gr-qc/9402012]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1016/0370-1573(94)00111-F}{\detokenize{10.1016/0370-1573(94)00111-F}}}.

\bibitem[Wallner(1982)]{Wallner:1981jf}
Wallner, R.P.
\newblock {Notes on Gauge Theory and Gravitation}.
\newblock {\em Acta Phys. Austriaca} {\bf 1982}, {\em 54},~165.

\bibitem[Jiménez \em{et~al.}(2018)Jiménez, Heisenberg, and
  Koivisto]{BeltranJimenez:2017tkd}
Jiménez, J.B.; Heisenberg, L.; Koivisto, T.
\newblock {Coincident General Relativity}.
\newblock {\em Phys. Rev.} {\bf 2018}, {\em D98},~044048,
  \href{http://xxx.lanl.gov/abs/1710.03116}{{\normalfont
  [arXiv:gr-qc/1710.03116]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1103/PhysRevD.98.044048}{\detokenize{10.1103/PhysRevD.98.044048}}}.

\bibitem[Beltrán~Jiménez \em{et~al.}(2018)Beltrán~Jiménez, Heisenberg, and
  Koivisto]{BeltranJimenez:2018vdo}
Beltrán~Jiménez, J.; Heisenberg, L.; Koivisto, T.S.
\newblock {Teleparallel Palatini theories}.
\newblock {\em JCAP} {\bf 2018}, {\em 1808},~039,
  \href{http://xxx.lanl.gov/abs/1803.10185}{{\normalfont
  [arXiv:gr-qc/1803.10185]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1088/1475-7516/2018/08/039}{\detokenize{10.1088/1475-7516/2018/08/039}}}.

\bibitem[Fontanini \em{et~al.}(2019)Fontanini, Huguet, and
  Le~Delliou]{Fontanini:2018krt}
Fontanini, M.; Huguet, E.; Le~Delliou, M.
\newblock {Teleparallel gravity equivalent of general relativity as a gauge
  theory: Translation or Cartan connection?}
\newblock {\em Phys. Rev.} {\bf 2019}, {\em D99},~064006,
  \href{http://xxx.lanl.gov/abs/1811.03810}{{\normalfont
  [arXiv:gr-qc/1811.03810]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1103/PhysRevD.99.064006}{\detokenize{10.1103/PhysRevD.99.064006}}}.

\bibitem[Pereira and Obukhov(2019)]{Pereira:2019}
Pereira, J.G.; Obukhov, Y.N.
\newblock {Gauge Structure of Teleparallel Gravity}.
\newblock {\em Universe} {\bf 2019}, {\em to appear}.

\bibitem[Jiménez \em{et~al.}(2019)Jiménez, Heisenberg, and
  Koivisto]{Jimenez:2019yyx}
Jiménez, J.B.; Heisenberg, L.; Koivisto, T.S.
\newblock {The canonical frame of purified gravity} {\bf 2019}.
\newblock  \href{http://xxx.lanl.gov/abs/1903.12072}{{\normalfont
  [arXiv:gr-qc/1903.12072]}}.

\bibitem[Koivisto(2018)]{Koivisto:2018aip}
Koivisto, T.
\newblock {An integrable geometrical foundation of gravity}.
\newblock {\em Int. J. Geom. Meth. Mod. Phys.} {\bf 2018}, {\em 15},~1840006,
  \href{http://xxx.lanl.gov/abs/1802.00650}{{\normalfont
  [arXiv:gr-qc/1802.00650]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1142/S0219887818400066}{\detokenize{10.1142/S0219887818400066}}}.

\bibitem[Wise(2010)]{Wise:2006sm}
Wise, D.K.
\newblock {MacDowell-Mansouri gravity and Cartan geometry}.
\newblock {\em Class. Quant. Grav.} {\bf 2010}, {\em 27},~155010,
  \href{http://xxx.lanl.gov/abs/gr-qc/0611154}{{\normalfont
  [arXiv:gr-qc/gr-qc/0611154]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1088/0264-9381/27/15/155010}{\detokenize{10.1088/0264-9381/27/15/155010}}}.

\bibitem[Gielen and Wise(2013)]{Gielen:2012fz}
Gielen, S.; Wise, D.K.
\newblock {Lifting General Relativity to Observer Space}.
\newblock {\em J. Math. Phys.} {\bf 2013}, {\em 54},~052501,
  \href{http://xxx.lanl.gov/abs/1210.0019}{{\normalfont
  [arXiv:gr-qc/1210.0019]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1063/1.4802878}{\detokenize{10.1063/1.4802878}}}.

\bibitem[Westman and Zlosnik(2015)]{Westman:2014yca}
Westman, H.F.; Zlosnik, T.G.
\newblock {An introduction to the physics of Cartan gravity}.
\newblock {\em Annals Phys.} {\bf 2015}, {\em 361},~330--376,
  \href{http://xxx.lanl.gov/abs/1411.1679}{{\normalfont
  [arXiv:gr-qc/1411.1679]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1016/j.aop.2015.06.013}{\detokenize{10.1016/j.aop.2015.06.013}}}.

\bibitem[Hohmann(2016)]{Hohmann:2015pva}
Hohmann, M.
\newblock {Spacetime and observer space symmetries in the language of Cartan
  geometry}.
\newblock {\em J. Math. Phys.} {\bf 2016}, {\em 57},~082502,
  \href{http://xxx.lanl.gov/abs/1505.07809}{{\normalfont
  [arXiv:math-ph/1505.07809]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1063/1.4961152}{\detokenize{10.1063/1.4961152}}}.

\bibitem[Złośnik and Westman(2017)]{Zlosnik:2016fit}
Złośnik, T.G.; Westman, H.F.
\newblock {A first-order approach to conformal gravity}.
\newblock {\em Class. Quant. Grav.} {\bf 2017}, {\em 34},~245001,
  \href{http://xxx.lanl.gov/abs/1601.00567}{{\normalfont
  [arXiv:gr-qc/1601.00567]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1088/1361-6382/aa944f}{\detokenize{10.1088/1361-6382/aa944f}}}.

\bibitem[Capovilla \em{et~al.}(1991)Capovilla, Jacobson, Dell, and
  Mason]{Capovilla:1991qb}
Capovilla, R.; Jacobson, T.; Dell, J.; Mason, L.J.
\newblock {Selfdual two forms and gravity}.
\newblock {\em Class. Quant. Grav.} {\bf 1991}, {\em 8},~41--57.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1088/0264-9381/8/1/009}{\detokenize{10.1088/0264-9381/8/1/009}}}.

\bibitem[Nesti and Percacci(2008)]{Nesti:2007ka}
Nesti, F.; Percacci, R.
\newblock {Graviweak Unification}.
\newblock {\em J. Phys.} {\bf 2008}, {\em A41},~075405,
  \href{http://xxx.lanl.gov/abs/0706.3307}{{\normalfont
  [arXiv:hep-th/0706.3307]}}.
\newblock
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1088/1751-8113/41/7/075405}{\detokenize{10.1088/1751-8113/41/7/075405}}}.

\bibitem[Weyl(1918)]{Weyl:1918ib}
Weyl, H.
\newblock {Gravitation and electricity}.
\newblock {\em Sitzungsber. Preuss. Akad. Wiss. Berlin (Math. Phys.)} {\bf
  1918}, {\em 1918},~465.
\newblock [,24(1918)].

\bibitem[Weyl(1929)]{Weyl:1929fm}
Weyl, H.
\newblock {Electron and Gravitation. 1. (In German)}.
\newblock {\em Z. Phys.} {\bf 1929}, {\em 56},~330--352.
\newblock [,121(1929)],
  doi:{\changeurlcolor{black}\href{https://doi.org/10.1007/BF01339504}{\detokenize{10.1007/BF01339504}}}.

\bibitem[Wheeler(2018)]{Wheeler:2018qxx}
Wheeler, J.T.
\newblock {General relativity as a biconformal gauge theory} {\bf 2018}.
\newblock  \href{http://xxx.lanl.gov/abs/1808.07083}{{\normalfont
  [arXiv:gr-qc/1808.07083]}}.

\end{thebibliography}

\documentclass[preprints,article,accept,moreauthors,pdflatex]{Definitions/mdpi}

\usepackage{tipa}
\usepackage{amsmath,latexsym,mathrsfs,amssymb}

\newcommand{\lp}{\left(}
\newcommand{\rp}{\right)}
\newcommand{\lb}{\left[}
\newcommand{\rb}{\right]}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\ba}{\begin{eqnarray}}
\newcommand{\ea}{\end{eqnarray}}

\newcommand{\diff}{{{\rm d}}}
\newcommand{\Diff}{{{\rm D}}}

\newcommand{\bbe}{\boldsymbol{\rm e}}
\newcommand{\e}{\mathrm{e}}
\newcommand{\ie}{\text{\textschwa}}
\newcommand{\bae}{\boldsymbol{\text{\ae}}}
\newcommand{\bbie}{\boldsymbol{\textsf{\textschwa}}}

\newcommand{\bepsilon}{\boldsymbol{\epsilon}}

\newcommand{\bM}{\boldsymbol{\mathrm{M}}}

\newcommand{\bz}{\boldsymbol{\mathrm{z}}}

\newcommand{\bomega}{\boldsymbol{{\omega}}}

\newcommand{\bT}{\mathbf{T}}

\newcommand{\bF}{\mathbf{F}}
\newcommand{\bR}{\mathbf{R}}
\newcommand{\bx}{\mathbf{x}}
\newcommand{\by}{\mathbf{y}}
\newcommand{\bt}{\mathbf{t}}
\newcommand{\bs}{\mathbf{s}}
\newcommand{\bS}{\mathbf{S}}

\firstpage{1} 
\pubvolume{xx}
\issuenum{1}
\articlenumber{5}
\pubyear{2019}
\copyrightyear{2019}
\history{Received: date; Accepted: date; Published: date}
\pdfoutput=1

\Title{The General Linear Cartan Khronon}

\Author{Tomi Koivisto $^{1,2,3}$\orcidA{}*, Manuel Hohmann$^{2}$, Tom Z\l{}o\'s{}nik$^{4}$}

\address{%
$^{1}$ \quad Nordita, KTH Royal Institute of Technology and Stockholm University, Roslagstullsbacken 23, SE-10691 Stockholm, Sweden; tomi.koivisto@nordita.org\\
$^{2}$ \quad Laboratory of Theoretical Physics, Institute of Physics, University of Tartu, W. Ostwaldi 1, 50411 Tartu, Estonia \\
$^{3}$ \quad National Institute of Chemical Physics and Biophysics, R\"avala pst. 10, 10143 Tallinn, Estonia \\
$^{4}$ \quad CEICO, Institute of Physics of the Czech Academy of Sciences, Na Slovance 1999/2, 182 21, Prague}
\corres{Correspondence: tomik@astro.uio.no}

\abstract{A Cartan geometry of the General Linear symmetry is formulated by dividing out the displacements from the group. The resulting action is quadratic in curvature, polynomial in all the (minimal) variables, and describes an observer space that, in the symmetry-broken phase, reproduces the predictions of General Relativity in the presence of dark matter.}

\keyword{Conformal and Metric-Affine Gauge Theories of Gravitation, Cartan Geometry.}

\begin{document}

All energy and momentum source gravity. Alone due to this observation one naturally entertains the idea that ''gravity is that field which corresponds to a gauge invariance with respect to displacement transformations'' \cite{Feynman:1996kb}. 
%by energy and momentum, it is the field corresponding to gauge invariance with respect to displacement transformations \cite{Feynman:1996kb} i.e. translations. 
It is often considered that torsion $T^a$ is the field strength of displacements. However, since the coframe one-form $\e^a=t^a + \mathcal{D}x^a$ \cite{Grignani:1991nj} is the displacement gauge potential $t^a$ only up to the covariant derivative, $\mathcal{D}=\diff + \bomega$ where $\bomega$ is the Lorenz connection, of a coordinate-scalar $x^a$, the torsion \cite{cartan1986manifolds} 
\be
T^a = \mathcal{D}\e^a = \mathcal{D}t^a - R_b{}^a x^b\,,
\ee  
is equal to the displacement field strength, $T^a = \mathcal{D}t^a$, only if the curvature $R_a{}^b=\mathcal{D}\omega_a{}^b$ vanishes i.e. only if teleparallelism, $R_a{}^b=0$, is assumed. In this sense, teleparallel gravity \cite{Aldrovandi:2013wha} can be regarded as a gauge theory of displacements, but so can just as well the more conventional
description of gravity without torsion, since if $T^a=0$, the displacement gauge field strength $\mathcal{D}t^a=R_b{}^a x^b$ is in turn directly proportional to the curvature. In fact, this form is manifest in the recently discovered action that is polynomial in all the variables and quadratic in curvature,
\be
L = -\frac{1}{2}\lp\frac{1}{8}x^2\epsilon^{ac}{}_{bd} + ix^a x^c\eta_{bd}\rp R_{a}{}^{b}\wedge R_{c}{}^{d}\,, \label{cc}
\ee  
but reproduces the predictions of General Relativity in the presence of geometrical dark matter and is possibly capable of extending the description of spacetime into a pre-geometric phase, without invoking any (co)frame field \cite{Zlosnik:2018qvg} (the four-form (\ref{cc}) is considered as a function of only $\omega_a{}^b$ and $x^a$). These are fresh insights, but follow closely Cartan's original developments \cite{cartan1986manifolds} , where attention was directed to the case wherein the displacement gauge potential vanishes, $t^a=0$. In such a case the field $x^a$ is known as the Cartan radius vector \cite{Hehl:1994ue}. In the context of the new theory \cite{Zlosnik:2018qvg} we might call it the {\it Cartan Khronon} since it is plays the role of a symmetry-breaking field that determines the direction of time (which only exists in the geometric, symmetry-broken phase).

Besides a Cartan-geometric, also a canonical gauge theory of displacements \cite{Wallner:1981jf} has been taken into reconsideration in the current literature \cite{BeltranJimenez:2017tkd,BeltranJimenez:2018vdo}. A displacement transformation can be seen, from the passive perspective, as a general coordinate transformation. Obviously, such a transformation is included in the GL (General Linear) transformations\footnote{This is contrary to what was implied in Refs. \cite{Fontanini:2018krt,Pereira:2019}. We agree with Ref. \cite{Pereira:2019} that the principal bundle of the displacement group is not a vector bundle, since in gauge theories the fibres are actually torsors instead of groups (which though does not change the conclusion of Ref. \cite{Fontanini:2018krt} that such a displacement bundle is topologically trivial).}.  In a GL gauge connection, the displacement component is precisely that which has neither curvature nor torsion\footnote{Consider an affine connection $\Gamma^\alpha{}_{\mu\nu}$ in the tensor representation. If without curvature, it will be given by a GL transformation $\Lambda^\alpha{}_\beta$ of the zero connection as $\Gamma^\alpha{}_{\mu\nu}=\Lambda^\alpha_\beta\partial_\mu(\Lambda^{-1}){}^\beta{}_\nu$, and if one further assumes the vanishing of torsion $T^\alpha{}_{\mu\nu}=2\Gamma^\alpha{}_{[\mu\nu]}=0$, it follows further that  $\Lambda^\alpha{}_\beta=\partial_\beta \xi^\alpha$, which is nothing but the Jacobian of a coordinate transformation.} \cite{BeltranJimenez:2017tkd}. The reformulation of Einstein's theory in such a geometrical setting which is ''purified'' from all non-integrable elements allows the localisation of the gravitational energy and the path integral formalism that does not suffer from the necessity of boundary terms \cite{Jimenez:2019yyx}. The unique property of the gravitational interaction that it can be always locally eliminated, is reflected in the integrability of the connection \cite{Koivisto:2018aip}. 

Thus, both the canonical and the Cartan-geometric approaches to gauge theory of displacements suggest that $\mathcal{D}t^a=0$, and the physical interpretation of this is the essence of General Relativity, the equivalence of gravity and inertia. In this note, we shall find some further connections between between purified gravity and Cartan geometry \cite{cartan1986manifolds,Wise:2006sm,Gielen:2012fz,Westman:2014yca,Hohmann:2015pva}. The purpose of this note is to formulate a gauge theory of displacements in terms of a GL-covariant (exterior) derivative $\Diff$, and two symmetry-breaking fields $\bx$ and $\by$. The spacetime metric will then be an emergent structure that only exists
in the symmetry-broken phase. 

%In  that the corresponding field is pure gauge, i.e. $\mathcal{D}t^a=0$
%From a more mathematical perspective, the theory of principal bundles forces the canonical gauge theory of displacements to have null field strength because the tangent spaces areisomorphic 
%This was first realised in the theory of coincident GR, which is indeed is {\it the} gauge theory of translations understood from the passive perspective as general coordinate transformations: iThe canonical gauge theory was formulated in terms of this integrable connection, and its non-metricity $\nabla_\mu g_{\alpha\beta}$ with respect to the metric $g_{\mu\nu}$. 

In the conventional metric-affine gauge theories of gravitation \cite{Hehl:1994ue}, besides invoking the metric as an additional field, one gauges both the GL symmetry and displacements. 
In the language of Cartan geometry, the model is $G/H$, where $G = \mathrm{GA}(4, \mathbb{R}) = \mathrm{GL}(4, \mathbb{R}) \ltimes \mathbb{R}^4$, sometimes called the general affine group, is an inhomogenenisation of the
$H = \mathrm{GL}(4, \mathbb{R})$. However, this is superfluous since the displacements are already included in the GL as pointed out above. Actually, there are two (coupled) sets of displacements. This can be made apparent in the spinor representation in terms of the Dirac matrices $\gamma_a$ which satisfy
the Clifford algebra $\{\gamma_a, \gamma_b \} = -2\eta_{ab}$. We have the left and the right displacements, generated by 
\be
\overset{+}{\gamma}_a=\frac{1}{2}\lp 1+\gamma_5\rp\gamma_a\,, \quad \overset{-}{\gamma}{}^a=\frac{1}{2}\lp 1- \gamma_5\rp\gamma^a\,, \label{trans}
\ee
which form two Abelian subgroups. We write the full GL connection and its curvature as
\be 
\Diff = \diff +  t^a\overset{+}{\gamma}_a + s_a\overset{-}{\gamma}{}^a +  \frac{1}{2}\omega_{a}{}^{b}i\sigma^a{}_{b} + p\frac{-\gamma_5}{2} + q\,, \quad 
\bF = \tilde{T}^a\overset{+}{\gamma}_a + \tilde{S}_a\overset{-}{\gamma}{}^a +  \frac{1}{2}R_{a}{}^{b}i\sigma^{a}{}_{b} - \frac{1}{2}P\gamma_5 + Q\,, \label{second}
\ee 
respectively. In this basis we may recognise the connection of conformal gauge theory \cite{Zlosnik:2016fit} plus the trivial gauge field $q$.
Explicitly, the field strength components in (\ref{second}) are
\begin{subequations}
\ba
\tilde{T}^a & = & \diff t^a - \omega_b{}^a\wedge t^b-p\wedge t^a\,, \\
\tilde{S}_a & = & \diff s_a + \omega_a{}^b\wedge s_b + p\wedge s_a\,, \\
R_a{}^b & = &  \diff\omega_a{}^b - \omega_a{}^c\wedge\omega_c{}^b - 2t^{[b}\wedge s_{a]}\,, \\
P & = & \diff p -t^a\wedge s_a\,, \\
Q & = & \diff q\,.
\ea
\end{subequations}
The nonvanishing algebra includes the Poincar\'e subalgebras
\be
[\sigma_{ab},\sigma_{cd}]=2i\lp \eta_{c[a}\sigma_{b]d}-\eta_{d[a}\sigma_{b]c}\rp\,, \,\,\,
[\overset{\pm}{\gamma}_a,\sigma_{bc}]=2i\eta_{a[c}\overset{\pm}{\gamma}_{b]}\,,  \label{poincare}
\ee
and
\be
[\overset{+}{\gamma}_a,\overset{-}{\gamma}_b] = 2\lp\eta_{ab}\frac{-\gamma_5}{2}-i\sigma_{ab}\rp\,,  \,\,\, 
[\overset{\pm}{\gamma}_a,\frac{-\gamma_5}{2}] = \pm \overset{\pm}{\gamma}_a\,, \quad\quad \label{rest}
\ee
is the rest of the algebra. 
Note in particular the presence of a 8-dimensional subgroup $H$ generated by $\sigma^a{}_b, \gamma_5$ and the central element $1$ of the Clifford algebra, which together with $G =$ GL may define a Cartan gauge theory modeled on $G/H$. 
More conventional 4-dimensional model geometry is achieved instead by also including one set of displacements into the $H$ that is then 12-dimensional. 
From the algebra relations we can determine the Cartan-Killing form\footnote{Or alternatively, from the traces of the matrix products,
noting that for GL(n) matrices $\bM_1$ and $\bM_2$ we have $\langle \bM_1\bM_2\rangle = \text{Tr}(\bM_1\bM_2)-\text{Tr}(\bM_1)\text{Tr}(\bM_2)/n$. To display compactly the algebra (\ref{poincare},\ref{rest},\ref{killing}) we have raised and lowered the indices of the generators with the metric $\eta_{ab}$, though elsewhere in this note we avoid that.}. Its nonzero components dictate that
\be
\langle \sigma^{ab}\sigma_{cd}  \rangle = 4\delta^{[a}_c\delta^{b]}_d\,, \quad \langle \overset{+}{\gamma}_a \overset{-}{\gamma}_b \rangle = \langle \overset{-}{\gamma}_a \overset{+}{\gamma}_b \rangle  = -2\eta_{ab}\,, \quad \langle \gamma_5\gamma_5 \rangle = 4\,. \label{killing}
\ee 
Thus we have a group-invariant definition of an inner product. 

To incorporate spinors (and quantum mechanics), it is necessary to consider the group GL(4,$\mathbb{C}$) over complex numbers, since the GL(4,$\mathbb{R})$ over real numbers does not have finite-dimensional spinors. It is convenient then to take $\eta_{ab}=-\delta_{ab}$ (and thus, Euclidean gamma matrices).
 Though GL(4,$\mathbb{C}$) has the real dimension 32, we would like to point out in passing that actually, there may nothing extra in the connection (\ref{second}) if one eventually is aiming at a unification: 
like in Plebanski's and Ashtekar's gravity \cite{Capovilla:1991qb}, the spin connection is the self-dual part of $\omega_{a}{}^{b}$; like in the Cartan Khronon theory \cite{Zlosnik:2018qvg}, the real component of the anti-self dual 
part of $\omega_{a}{}^{b}$ could be essentially the triad; as in the graviweak theory \cite{Nesti:2007ka} the imaginary component of the anti-self dual could be the SU$_\text{L}$(2); and along the lines of Weyl's gauge theories \cite{Weyl:1918ib,Weyl:1929fm}, the real part of the dilaton $p$ 
could gauge the scale invariance and the imaginary part the U$_\text{Y}$(1). Since in the spinorial representation (\ref{poincare},\ref{rest}) it becomes apparent that the SL(4,$\mathbb{C}$) double-covers the SO(6,$\mathbb{C}$), we have dubbed this the Conformal Affine Theory. An interesting property is that the $-\gamma_5/2$ has the role of the generator of dilatons in spacetime geometry. This hints at the connection
between spatial length scales and particle mass scales, since in quantum field theory the appearance of mass terms for matter fields occur via left-right couplings. 
In this note, however, our focus is on gravity and the displacements (\ref{trans}). 

We now begin the Cartan construction with two spacetime scalars $\bx=x^a\overset{+}{\gamma}_a$ and $\by=y_a\overset{-}{\gamma}{}^a$,
\begin{subequations}
\ba
\bbe & =&  \bt + \Diff \bx = \lb t^a + \lp \diff -p\rp x^a - \omega_b{}^a x^b\rb\overset{+}{\gamma}_a  + s_a x^a \gamma_5 + 2 s_a x^b i\sigma^{a}{}_{b} \,, \\
\bbie & = & \bs + \Diff \by =  \lb s_a + \lp \diff + p\rp y_a + \omega_a{}^b y_b\rb\overset{-}{\gamma}{}^a  - t^a y_a \gamma_5 + 2 t^b y_a i\sigma^{a}{}_{b} \,.
\ea
\end{subequations}
Only in the gauge $\bt=0$ the $\bbie$ becomes a conventional frame field for any $\by$, and only in the gauge $\bs=0$ the $\bbe$ becomes a conventional coframe field for any $\bx$. The availability of these
gauge choices requires that $\tilde{\bT}=0$ and $\tilde{\bS}=0$, respectively. The torsion and cotorsion are
\begin{subequations}
\ba
\bT= \Diff \bbe & =& \lp \tilde{T}^a - R_b{}^a x^b - Px^a\rp\overset{+}{\gamma}_a + \lp s_a\wedge t^a + \tilde{S}_a x^a\rp \gamma_5 - 2\lp  s_a\wedge t^b + \tilde{S}_ax^b\rp i\sigma^{a}{}_{b} \,, \\
\bS= \Diff \bbie & =& \lp \tilde{S}_a + R_a{}^b y_b + Py_a\rp\overset{-}{\gamma}{}^a - \lp s_a\wedge t^a + \tilde{T}^a y_a\rp \gamma_5 + 2\lp  s_a\wedge t^b + \tilde{T}^b y_a\rp i\sigma^{a}{}_{b}  \,.
\ea
\end{subequations}
Consider the gauge transformation
\be
\bepsilon = \overset{+}{\epsilon}{}^a\overset{+}{\gamma}_a + \overset{-}{\epsilon}{}_a\overset{-}{\gamma}{}^a + \frac{1}{2}\epsilon_{a}{}^{b}i\sigma^{a}{}_{b} - \frac{1}{2}\epsilon\gamma_5+ \varepsilon I\,,
\ee
where $\eta^{ca}\epsilon_c{}^b=\eta^{c[a}\epsilon_c{}^{b]}$.
The gauge potentials transform as 
\begin{subequations}
\ba
\delta_{\bepsilon} t^a & = & -\lp \diff - p\rp\overset{+}{\epsilon}{}^a + \omega_b{}^a \overset{+}{\epsilon}{}^b - t^b\epsilon_b{}^a - t^a\epsilon\,, \\
\delta_{\bepsilon} s_a & = &  -\lp \diff + p\rp\overset{-}{\epsilon}{}_a - \omega_a{}^b \overset{-}{\epsilon}{}_b + s_b\epsilon_a{}^b + s_a\epsilon\,, \\
\delta_{\bepsilon} \omega_{a}{}^{b} & = & -\diff\epsilon_{a}{}^{b} - 2\omega_c{}^{[a}\epsilon_{b]}{}^{c}  -4\lp s_{[a}\overset{+}{\epsilon}{}^{b]} - t^{[a}\overset{-}{\epsilon}{}_{b]}\rp \,, \\
\delta_{\bepsilon} p & = & -\diff\epsilon-2\lp s_a \overset{+}{\epsilon}{}^a - t^a\overset{-}{\epsilon}{}_a\rp \,, \\
\delta_{\bepsilon} q & = & -\diff\varepsilon\,,
\ea
\end{subequations}
and the two GL-charged coordinate-scalars transform as
\begin{subequations}
\ba
\delta_{\bepsilon} \bx & = & \lp \overset{+}{\epsilon}{}^a - \epsilon_b{}^a x^b - \epsilon x^a\rp\overset{+}{\gamma}_a + \overset{-}{\epsilon}{}_a x^a\gamma_5 + 2\overset{-}{\epsilon}{}_a x^b i\sigma^{a}{}_{b}  \,, \\
\delta_{\bepsilon} \by & = & \lp \overset{-}{\epsilon}{}_a + \epsilon_a{}^b y_b + \epsilon y_a\rp\overset{-}{\gamma}{}^a - \overset{+}{\epsilon}{}^a y_a\gamma_5 + 2\overset{+}{\epsilon}{}^a y_b i\sigma^b{}_a\,.
\ea
\end{subequations}
%Because the Cartan geometry is not reductive \cite{Wise:2006sm}, we need to consider some restrictions. 
We can check the transformations of the components $\e^a$ and $\ie_a$ to verify that
\begin{itemize}
\item If we fix $\overset{-}{\bepsilon}$ such that $\bs=0$, the cotetrad transforms as $\delta_{\bepsilon}\e^a = -\epsilon_b{}^a\e^b + \epsilon\e^a$, in particular $\delta_{\overset{+}{\bepsilon}}\bbe=0$. 
\item If we fix $\overset{+}{\bepsilon}$ such that $\bt=0$, the tetrad transforms as $\delta_{\bepsilon}\ie_a = \epsilon_a{}^b\ie_b - \epsilon\ie_a$, in particular $\delta_{\overset{-}{\bepsilon}}\bbie=0$. 
\end{itemize}
Imposing both would essentially leave us with the centrally extended Weyl group (generated by  the central element $1$, $\sigma^a{}_b$ and $\gamma_5$) properly gauged, and imposing one of the above conditions would leave us with the centrally extended inhomogeneous Weyl group (which has either the $\overset{-}{\gamma}{}^a$ or the $\overset{+}{\gamma}_a$ as additional generators) properly gauged. It is because the Cartan geometry is not reductive \cite{Wise:2006sm}, that we need to impose some restrictions on the solutions.

In either context\footnote{There would be interesting alternative ways to proceed \cite{Koivisto:2018aip,Wheeler:2018qxx}, but we need not be concerned with them in this note.} we can embed the Cartan Khronon into the GL gauge theory. 
%Using the scalar Higgs field $\bX = \frac{1}{2}x^a\lp \overset{+}{\gamma}_a + \overset{-}{\gamma}_a\rp$ and 
Using the Cartan-Killing form (\ref{killing}), one can write an invariant Lagrangian four-form involving the GL curvature and the Cartan 
Khronon as follows:
%\footnote{This could be interpreted also as the sum of the kinetic terms for the gauge fields and for the symmetry-breaking field, since equivalently $4iL =  \frac{1}{4}\langle \bx\bF\wedge\by\bF\rangle + (\Diff^2x^a\wedge \Diff^2 y^b)\langle \overset{+}{\gamma}_a \overset{-}{\gamma}_b\rangle$.}
\be \label{l2}
L = -\frac{i}{8}\lp \langle \bx\bF\wedge\by\bF\rangle + \frac{1}{2} \langle[\bx,\by]\bF\wedge\bF\rangle  
 - \frac{1}{4}\langle \bx\by\rangle\langle \bF\wedge\bF\rangle\rp\,. 
%\,\,\, \rightarrow \,\,\, \frac{1}{2}\lp\frac{1}{8}x^2\epsilon_{abcd}-ix_ax_c\eta_{bd}\rp R^{ab}\wedge R^{cd}\,.
%L = -i\langle\frac{1}{4}[\bx,\by]\bF\wedge\bF  + \Diff^2 \bx \wedge \Diff^2 \by\rangle \,\,\, \rightarrow \,\,\, \frac{1}{2}\lp\frac{1}{8}x^2\epsilon_{abcd}-ix_ax_c\eta_{bd}\rp R^{ab}\wedge R^{cd}\,.
%L = \frac{1}{4}\epsilon_{ABCDEF} x^{AG}y^B{}_G F^{CD}\wedge F^{EF} - 2i\langle \Diff^2 \bx, \wedge \Diff^2 \by\rangle \,\,\, \rightarrow \,\,\, \frac{1}{2}\lp\frac{1}{8}x^2\epsilon_{abcd}-ix_ax_c\eta_{bd}\rp R^{ab}\wedge R^{cd}\,.
\ee
We have chosen the coefficients of the invariants such that by setting $\bF \rightarrow \bR$ and $y_a \rightarrow \eta_{ab}x^b$ (\ref{l2}) reduces precisely to\footnote{It seems evident that the 
equations of motion for $s_a$ and $t^a$ are trivially satisfied by $s_a=t^a=0$. Note that the dilation invariance is retained.} (\ref{cc}). 
We have not introduced a metric by hand, but its components $g_{\mu\nu}$ will correspond to $-{\footnotesize{{\frac{1}{2}}}}\langle \bbie, \bbe\rangle  \rightarrow \eta_{ab}\mathcal{D}x^a\otimes\mathcal{D} x^b = g_{\mu\nu}\diff x^\mu\otimes\diff x^\nu$. 
The compatible connection is given by $-2{\Gamma}^\alpha{}_{\mu\nu} =  \ie_a{}^\alpha\mathcal{D}_\mu\e^a{}_\nu = -\e^a{}_\nu\mathcal{D}_\mu\ie_a{}^\alpha$, whereas the connection of purified gravity is simply given by the same expression when $\mathcal{D} \rightarrow \diff$.
%By allowing for the Weyl curvature $\bF \rightarrow \bR + \bP$ and setting $y^a \rightarrow x^a$ in (\ref{l2}) we obtain
%\be
%L_{W} \rightarrow \frac{1}{2}\lp\frac{1}{8}x^2\epsilon_{abcd}-ix_ax_c\eta_{bd}\rp R^{ab}\wedge R^{cd} - ix^2P\wedge P\,,
%\ee
%a Weyl

Finally, we note that the GL Cartan Khronon can be considered as the single field $\bz = (\bx + i\by)/2$. This allows to replace (\ref{l2}) by 
\be \label{l3}
L =  \frac{1}{4}\lp  \langle \bz \bz^\ast \bF \wedge\bF\rangle - \langle \bz\bF\wedge\bz\bF\rangle  + \frac{1}{4}\langle \bz\bz\rangle\langle\bF\wedge \bF\rangle \rp\,. 
\ee 
If we also define $\bae = {\footnotesize{\frac{1}{2}}}(\bbe + i\bbie) = {\footnotesize{\frac{1}{2}}}(\bt + i\bs) + \Diff \bz$, the spacetime metric and connection will then correspond to the GL-invariant
objects $i\langle \bae \otimes \bae\rangle \rightarrow g_{\mu\nu}\diff x^\mu\otimes\diff x^\nu$ and  $i\langle \bae\otimes \Diff \otimes \bae\rangle \rightarrow \Gamma^\alpha{}_{\mu\nu}\partial_\alpha\otimes\diff x^\mu\otimes\diff x^\nu$, respectively, in the above mentioned limit wherein (\ref{l3}) reduces to the Lorentz
gauge theory (\ref{cc}). 

%Let us consider a different prescription. We may fix the combination $\overset{+}{\epsilon}{}^a i\gamma_5\gamma_a$ such that $t^a = s^a$, and $\epsilon$ such that $p=0$, and choose $y^a=x^a$.
%We are then left with the one independent set of translations by ${\epsilon}{}^a\gamma_a$ and can entertain solutions for which $\bbe = \bbie$. 
To conclude, we have formulated a Cartan-geometric, polynomial GL gauge theory wherein displacements are generated by an Abelian subgroup of the GL. 
The result, Eq.(\ref{l3}), extends General Relativity by eliminating the need for dark matter and by incorporating a symmetric, pre-geometric phase. Thus, we can combine the gratifying properties of a canonical displacement gauge theory and the realisation of an observer space \'ala Cartan Khronon. It remains to be seen how far we may proceed with unification from this basis. The challenges would be to assign the gauge fields of the particle interactions within the GL connection and to join the Cartan Khronon and the Higgs that was observed in the Large Hadron Collider into a single scalar. It could also be interesting to consider whether the GL gauge theory could address the cosmological problems of dark energy and the (inflationary) beginning of the universe. 

\acknowledgments{This note has benefitted from interesting discussions with Luca Marzola and many of the participants of the workshop {\it Teleparallel Universes in Salamanca}.  
The work was supported by the Estonian Research Council through the Personal Research Funding project PRG356 ``Gauge Gravity'' and by the European Regional Development Fund through the Center of Excellence TK133 ``The Dark Side of the Universe''.}

\bibliography{theory}

\end{document}